OpenStudy (anonymous):

The difference of x and y is 14. The value of x is 3 more than twice the value of y. Write two equations and graph to find the value of x. x = 25 x = 11 x = -17 x = 4

OpenStudy (anonymous):

Please help me !!!!!!

OpenStudy (compassionate):

x - y = 14 x = 3 + 2y

OpenStudy (compassionate):

Now solve for x

OpenStudy (anonymous):

Lol. I dont know how to. That is why i put it on here

OpenStudy (compassionate):

x - y = 14 x = 3 + 2y So we know x = 3 + 2y So plug x into the first equation

OpenStudy (anonymous):

Ill just guess Lol

OpenStudy (compassionate):

Don't just guess. Just plug x into the first equation. We know x is equal to 3 + 2y, so take x - y = 14 and just plug it in.

OpenStudy (anonymous):

It really doesnt make any sense

OpenStudy (compassionate):

Of course it makes sense. You're just not understanding it. What part are you having trouble with?

OpenStudy (anonymous):

\[x-y=14\]\[x=2y+3\] subsitutes the value of x \[(2y+3)-y=14\]\[y+3=14\]\[y=11\]